Levine Inc., which produces a single product, has prepared the following standard cost sheet for one unit of the product. Direct materials (9 pounds at $2.20 per pound) $19.80 Direct labor (1 hours at $11.00 per hour) $11.00 During the month of April, the company manufactures 310 units and incurs the following actual costs. Direct materials purchased and used (1,600 pounds) $3,680 Direct labor (330 hours) $3,564 Compute the total, price, and quantity variances for materials and labor.

Total materials variance $enter a dollar amount select an option

Materials price variance $enter a dollar amount select an option

Materials quantity variance $enter a dollar amount select an option

Total labor variance $enter a dollar amount select an option

Labor price variance $enter a dollar amount select an option

Labor quantity variance $enter a dollar amount select an option
Answer #1

1) Total material variance = (310*19.80-3680) = 2458 F

Material price variance = (2.20*1600-3680) = 160 U

Material quantity variance = (310*9-1600)*2.2 = 2618 F

Total labor variance = (310*11-3564) = 154 U

Labor rate variance = (11*330-3564) = 66 F

Labor quantity variance = (310-330)*11 = 220 U